# Towards Automatic Evaluation and High-Quality Pseudo-Parallel Dataset Construction for Audio Editing: A Human-in-the-Loop Method

### v1

@@ Abstract
Audio editing aims to manipulate audio content based on textual descriptions, supporting tasks such as adding, removing, or replacing audio events. Despite recent progress, the lack of high-quality benchmark datasets and comprehensive evaluation metrics remains a major challenge for both assessing audio editing quality and improving the task itself. In this work, we propose a novel approach for audio editing task by incorporating expert knowledge into both the evaluation and dataset construction processes: 1) First, we establish AuditScore, the first comprehensive dataset for subjective evaluation of audio editing, consisting of over 6,300 edited samples generated from 7 representative audio editing frameworks and 23 system configurations. Each sample is annotated by professional raters on three key aspects of audio editing quality: overall Quality, Relevance to editing intent, and Faithfulness to original features. 2) Based on this dataset, we train AuditEval, the first model designed for automatic MOS-style scoring tailored to audio editing tasks. AuditEval addresses the critical lack of objective evaluation metrics and the prohibitive cost of subjective assessment in this field. 3) We further leverage AuditEval to evaluate and filter a large amount of synthetically mixed editing pairs, constructing a high-quality pseudo-parallel dataset by selecting the most plausible samples. Objective experiments validate the effectiveness of our expert-informed filtering strategy in yielding higher-quality data, while also revealing the limitations of relying solely on objective metrics. The dataset, codes and tools can be found at: https://github.com/NKU-HLT/AuditEval.
Introduction
@@
Benefiting from the construction of Text-to-Audio (TTA) and Audio Caption (AC) datasets(Kim et al. 2019; Gemmeke et al. 2017; Drossos, Lipping, and Virtanen 2020; Mei et al. 2024; Piczak 2015; tza 2002), audio generation has made significant strides in recent years. Current TTA models—spanning diffusion-based frameworks(Huang et al. 2023; Liu et al. 2023, 2024b; Ghosal et al. 2023; Chen et al. 2024; Majumder et al. 2024; Hung et al. 2024; Xue et al. 2024) and autoregressive Transformer-based systems(Agostinelli et al. 2023; Kreuk et al. ; Copet et al. 2023)—are capable of producing high-quality, diverse, and text-aligned audio outputs based on user prompts. However, beyond generation, audio editing remains equally essential and presents more formidable challenges. Unlike generation, which involves creating audio from scratch, audio editing requires the system to modify existing audio content according to user-provided textual instructions or target descriptions. Unlike generation, which creates audio from scratch, editing modifies existing content based on textual instructions or target descriptions. It includes basic event manipulations (e.g., adding, removing, replacing) and more advanced tasks like super-resolution and style transfer. Audio editing poses three major challenges: maintaining audio quality, ensuring precise control aligned with user intent, and preserving the unedited portions of the original audio, rather than generating an entirely new sample.
@@
However, the development of audio editing has been significantly hindered by the lack of high-quality paired datasets—an issue also observed in the vision domain. Constructing supervised audio editing datasets is particularly difficult: unlike text-to-audio data, naturally occurring audio pairs that reflect specific editing operations are virtually nonexistent. Manual collection is often prohibitively expensive or infeasible—especially for scenarios like replacing gunshots in a crowd with thunder while preserving alignment. As a result, most existing methods adopt training-free approaches based on generative diffusion models(Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) and inversion techniques(Song, Meng, and Ermon 2020; Mokady et al. 2023; Wu and De la Torre 2023; Huberman-Spiegelglas, Kulikov, and Michaeli 2024). These methods repurpose the generation and cross-modal alignment capabilities of pretrained TTA models, steering generation toward desired edits by adjusting prompts or conditioning signals(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a; Tsai et al. 2024). However, they heavily depend on the underlying TTA model and often lack robustness due to the absence of task-specific supervision.
@@
As an alternative to training-free approaches, supervised frameworks have also been explored by training models on explicitly constructed editing datasets. Audit(Wang et al. 2023b) represents the first supervised audio editing framework, introducing a triplet-based dataset and training a dedicated latent diffusion model (LDM). To date, it remains the only task-specific paired dataset for audio editing. This dataset is created through audio mixing and combinatorial pairing strategies, leveraging the compositional nature of audio—unlike images, audio signals can be naturally combined to simulate coherent edits. A similar construction paradigm is adopted in InstructME(Han et al. 2023), which applies mixing-based strategies to generate supervision signals for music editing tasks. Meanwhile, MusicGen(Copet et al. 2023) and Instruct-MusicGen(Zhang et al. 2024a) explore autoregressive approaches based on discrete audio-text tokens, supporting tasks such as continuation, transformation, and editing. WavCraft(lia 2024) further investigates instruction-based editing by employing large language models (LLMs) to interpret textual commands and orchestrate external audio and speech models for modification.
@@
Despite recent progress, there is still no unified standard for evaluating audio editing tasks. The absence of a widely accepted, open-source benchmark specifically designed for editing has led to fragmented evaluation protocols and a lack of consistent objective metrics. Currently, the most commonly used metric is the CLAP Score, which measures similarity between the edited audio and the target prompt in the embedding space, leveraging the CLAP model(Elizalde et al. 2023) for cross-modal comparison without requiring ground-truth audio. To assess preservation of unedited regions, some works(Paissan et al. 2023; Zhang et al. 2024b; Jia et al. 2025; Liu et al. 2024a) use the original audio as a reference and compute similarity between the original and edited versions. Others(Xu et al. 2024; Zhang et al. 2024b) construct pseudo-paired datasets via audio mixing—following a strategy similar to Audit and InstructME—and evaluate outputs by comparing them to mixed signals using conventional TTA metrics such as Fréchet Distance (FD), Fréchet Audio Distance (FAD)(Roblek et al. 2019), and KL divergence. However, such evaluation paradigms—relying on synthetic mixtures as ground truth—often fall short. Since high-quality audio edits may vary significantly and are not strictly aligned with a single reference, these metrics can fail to reliably capture true editing quality or model controllability.
@@
As a result, subjective human evaluation remains the most reliable method for assessing audio editing quality. Current subjective evaluation protocols typically consider three aspects: overall audio Quality(Zhang et al. 2024b; Jia et al. 2025), Relevance to the editing intent(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a), and Faithfulness to the original (unedited) content(Paissan et al. 2023; Xu et al. 2024; Zhang et al. 2024b; Manor and Michaeli 2024; Jia et al. 2025; Liu et al. 2024a). These criteria provide an intuitive and comprehensive assessment of audio editing performance. Nevertheless, compared to pure generation tasks such as Text-to-Audio (TTA), Text-to-Speech (TTS), or Text-to-Music (TTM), audio editing is inherently more complex, requiring deeper understanding and recognition of nuanced content. This results in increased costs for annotator training, longer evaluation durations, and greater human labor requirements in practice. In light of these challenges, developing automatic evaluation methods and building a dedicated evaluation system for audio editing is both meaningful and valuable for advancing research in this field.
@@
In this paper, we aim to bridge the evaluation gap in audio editing by introducing AuditEval, a scalable automatic MOS-style assessment framework, and subsequently leveraging it to build a high-quality pseudo-parallel dataset. Our contributions are threefold:
@@
  • Subjective Audio Editing Evaluation Dataset Construction. We construct AuditScore, the first comprehensive benchmark for subjective audio editing evaluation, consisting of over 6,300 edited samples from 7+ representative frameworks and 23+ system variants. Each sample is annotated by professional raters on three key aspects: Quality, Relevance, and Faithfulness.
  • Automatic MOS Prediction. Based on AuditScore, we train AuditEval, the first model tailored for automatic MOS-style scoring in audio editing. It effectively mitigates the lack of objective metrics and the high cost of human evaluation.
@@
  • High-Quality Pseudo-Parallel Audio Editing Dataset Construction. We further utilize AuditEval to assess and filter a large volume of synthetically mixed editing pairs, selecting only the most plausible samples to construct a high-quality pseudo-parallel dataset. Objective experiments validate the effectiveness of this expert-informed filtering strategy in producing superior-quality data, while also highlighting the shortcomings of relying exclusively on objective metrics.
Related Work
Training-free Audio Editing Method
@@
Training-free audio editing methods exploit pre-trained TTA diffusion models to perform editing without the need for task-specific fine-tuning. These approaches typically begin by inverting the input audio into the noise space of the diffusion model—that is, mapping from $x_{0}$ (the observed audio) to $x_{T}$ (a noisy latent representation). The resulting latent $x_{T}$ preserves the semantic and structural features of the original audio while providing a flexible representation for guided generation. Editing is then achieved by denoising $x_{T}$ into a modified output $\hat{x}_{0}$, conditioned on user instructions or target descriptions. Depending on the specific inversion strategy used, training-free audio editing methods can be broadly categorized into three types: SDEdit-based, DDIM-based, and DDPM-based approaches.
@@
DDIM-based approaches perform deterministic inversion of the diffusion process by exploiting the approximate reversibility of the underlying ODE with sufficiently small step sizes (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020; Mokady et al. 2023). This allows for editing by first reversing the generation trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}T$ (Eq. 2, where $\alpha$ is a hyperparameter and $\varepsilon{\theta}$ denotes the pretrained latent diffusion model), and then regenerating the audio using a modified prompt (Eq. 3), thereby achieving semantically guided edits. Building upon DDIM inversion, several studies (Xu et al. 2024; Zhang et al. 2024b; Jia et al. 2025; Liu et al. 2024a) incorporate enhancements such as attention map manipulation and null-text optimization, demonstrating effective results in controllable audio editing.
@@
DDPM-based approaches include CycleDiffusion(Wu and De la Torre 2023) and Edit-Friendly DDPM(Huberman-Spiegelglas, Kulikov, and Michaeli 2024). This kind of method leverages the stochastic nature of the DDPM sampling process—specifically, the stochastic noise injected at each diffusion step—to preserve the characteristics of the original input. By adding noise to the original input and storing intermediate latent states at each step, the model can compute the corresponding stochastic noise components (Eq.4, $\mu_{t}\left(\mathrm{x}_{t}\right)$ is the calculated expectation of ${x}_{t}$, $\sigma_{t}$ is the hyperparameter), which are then reused during the denoising process under the modified prompt to retain original content. Edit-Friendly DDPM further improves by altering the noise injection mechanism so that the noise added at each timestep is independent, thereby enhancing editing controllability and performance(Huberman-Spiegelglas, Kulikov, and Michaeli 2024). ZETA Editing(Manor and Michaeli 2024) is the first known method to apply DDPM inversion to audio editing, demonstrating superior results compared to SDEdit- and DDIM-inversion-based methods.
@@
Supervised Audio Editing Models
@@
In contrast to training-free methods, supervised audio editing approaches rely on explicitly constructed editing datasets and can be broadly divided into non-autoregressive and autoregressive paradigms. Models such as Audit and InstructME employ latent diffusion models conditioned on audio-text multimodal inputs. They utilize specialized conditioning mechanisms and position-aware loss functions to enable end-to-end training on pseudo-labeled editing pairs or instrument transfer datasets, achieving effective editing performance (Wang et al. 2023b; Han et al. 2023). On the other hand, autoregressive approaches generate edited audio by conditioning on discrete audio-text representations of the original content and edit instructions as context. Representative models like MusicGen and Instruction-MusicGen adopt transformer-based decoders to sequentially synthesize edited audio (Copet et al. 2023; Zhang et al. 2024a). While these models benefit from strong language-audio alignment and high generation quality, they often struggle with temporal inconsistencies between the original and edited audio.
Automatic Evaluation for Speech, Music and Audio
@@
Automatic subjective evaluation plays a vital role in assessing AI-generated outputs by simulating human perception in a scalable and repeatable manner. Compared to purely objective metrics, it provides a more intuitive and comprehensive measure of perceptual quality. A well-designed evaluator not only aligns with human preferences but also guides the development of future generative models. In the speech domain, automatic Mean Opinion Score (MOS) prediction has been widely adopted to evaluate tasks such as Text-to-Speech (Cooper et al. 2022; Wang et al. 2023a, 2024), Text-to-Voice (Tang et al. 2024), and speech enhancement(Reddy, Gopal, and Cutler 2021; Mittag et al. 2021; Chen et al. 2022). These models have shown promising results and effectively complement traditional objective metrics.
@@
In contrast, automatic evaluation for audio and music remains more challenging due to the greater content diversity—from natural sound events to complex musical structures—and the inherently looser alignment between audio and textual descriptions compared to the tighter correspondence in speech-transcription pairs. To address these challenges, MusicEval(Liu et al. 2025) is the first known framework to provide automatic scoring for music generation. It constructs a high-quality evaluation dataset and trains scoring models to assess both overall quality and text relevance, achieving strong correlation with human judgments. Meanwhile, Meta’s Audiobox Aesthetics(Tjandra et al. 2025) proposes a unified framework for speech, music, and audio evaluation, defining four core aesthetic dimensions that collectively provide a comprehensive and structured standard for automatic audio assessment from an aesthetic perspective.
Data Collection
Basic Information of AuditScore

### v2

@@ Abstract
Audio editing aims to manipulate audio content based on textual descriptions, supporting tasks such as adding, removing, or replacing audio events. Despite recent progress, the lack of high-quality benchmark datasets and comprehensive evaluation metrics remains a major challenge for both assessing audio editing quality and improving the task itself. In this work, we propose a novel approach for audio editing task by incorporating expert knowledge into both the evaluation and dataset construction processes: 1) First, we establish AuditScore, the first comprehensive dataset for subjective evaluation of audio editing, consisting of over 6,300 edited samples generated from 7 representative audio editing frameworks and 23 system configurations. Each sample is annotated by professional raters on three key aspects of audio editing quality: overall Quality, Relevance to editing intent, and Faithfulness to original features. 2) Based on this dataset, we systematically propose AuditEval, a family of automatic MOS-style evaluators tailored for audio editing, covering both SSL-based and LLM-based approaches. It addresses the lack of effective objective metrics and the prohibitive cost of subjective evaluation in this field. 3) We further leverage AuditEval to evaluate and filter a large amount of synthetically mixed editing pairs, mining a high-quality pseudo-parallel subset by selecting the most plausible samples. Comprehensive experiments validate that our expert-informed filtering strategy effectively yields higher-quality data, while also exposing the limitations of traditional objective metrics and the advantages of AuditEval. The dataset, codes and tools can be found at: https://github.com/NKU-HLT/AuditEval.
1 Introduction
@@
Benefiting from the construction of Text-to-Audio (TTA) and Audio Caption (AC) datasets Kim et al. (2019); Gemmeke et al. (2017); Drossos et al. (2020); Mei et al. (2024), audio generation has made significant strides in recent years. Current TTA models, including diffusion-based frameworks Huang et al. (2023); Liu et al. (2023, 2024a); Ghosal et al. (2023); Chen et al. (2024); Majumder et al. (2024); Hung et al. (2024); Xue et al. (2024) and autoregressive transformer-based systems Agostinelli et al. (2023); Kreuk et al. ; Copet et al. (2023), are capable of producing high-quality, diverse, and text-aligned audio outputs based on textual user prompts. However, beyond generation, audio editing remains equally essential but has received comparatively less attention. Unlike generation, which creates audio from scratch, audio editing modifies existing audio content based on textual instructions or target descriptions. It includes basic event manipulations (e.g., adding, removing, replacing) and more advanced tasks like super-resolution and style transfer. Audio editing faces three main challenges: maintaining the overall audio quality, executing edits that accurately reflect user intent, and preserving the unedited portions rather than generating a completely new sample.
@@
Currently, the development of audio editing has been significantly hindered by the lack of high-quality paired datasets. This has also been observed in the image editing domain. Constructing supervised audio editing datasets is particularly difficult. Unlike TTA data, naturally occurring audio pairs that reflect specific editing operations are virtually nonexistent. Manual collection is often prohibitively expensive or infeasible, especially for scenarios like replacing gunshots in a crowd with thunder while preserving alignment. As a result, most existing methods adopt training-free approaches based on generative diffusion models Ho et al. (2020); Song et al. (2020) and inversion techniques Song et al. (2020); Mokady et al. (2023); Wu and De la Torre (2023); Huberman-Spiegelglas et al. (2024). These methods repurpose the generation and cross-modal alignment capabilities of pretrained TTA models, steering generation toward desired edits by adjusting prompts or conditioning signals Paissan et al. (2023); Xu et al. (2024); Zhang et al. (2024b); Manor and Michaeli (2024); Jia et al. (2025); Liu et al. (2024b); Tsai et al. (2024). However, they heavily depend on the underlying TTA model and often lack robustness due to the absence of task-specific supervision.
@@
As an alternative to training-free approaches, supervised frameworks have also been explored by training models on explicitly constructed editing datasets. Audit Wang et al. (2023) represents the first supervised audio editing framework, introducing a triplet-based dataset and training a dedicated latent diffusion model (LDM). To date, it remains the only task-specific paired dataset for audio editing. This dataset is created through audio mixing and combinatorial pairing strategies. It leverages the compositional nature of audio, which unlike images, can be naturally mixed to simulate coherent edits. A similar construction paradigm is adopted in InstructME Han et al. (2023), which applies mixing-based strategies to generate supervision signals for music editing tasks. Meanwhile, MusicGen Copet et al. (2023) and Instruct-MusicGen Zhang et al. (2024a) explore autoregressive approaches based on discrete audio-text tokens, supporting tasks such as continuation, transformation, and editing. WavCraft Liang et al. further investigates instruction-based editing by employing large language models (LLMs) to interpret textual commands and call external audio and speech models for modification.
@@
Despite these progress, there is still no unified standard for evaluating audio editing tasks. The absence of a widely accepted, open-source benchmark specifically designed for editing has led to fragmented evaluation protocols and a lack of reliable objective metrics. Currently, the most commonly used metric is the CLAP_Score, which measures similarity between the edited audio and the target prompt in the embedding space, leveraging the CLAP model Elizalde et al. (2023) for cross-modal comparison without requiring ground-truth audio. To assess preservation of unedited regions, some works Paissan et al. (2023); Zhang et al. (2024b); Jia et al. (2025); Liu et al. (2024b) use the original audio as a reference and compute similarity between the original and edited versions. Others Xu et al. (2024); Zhang et al. (2024b) construct pseudo-paired datasets via audio mixing, following a strategy similar to that of Audit, and evaluate the generated outputs by comparing them with the mixed signals using conventional TTA metrics such as Fréchet Distance (FD), Fréchet Audio Distance (FAD) Roblek et al. (2019) and KL divergence. However, such evaluation paradigms, which rely on synthetic mixtures as ground truth, often fall short. Because high-quality audio edits can vary substantially and are not necessarily aligned with a single reference.
@@
As a result, subjective human evaluation remains the most reliable method for assessing audio editing quality. Current subjective evaluation protocols typically consider three aspects: overall audio Quality Zhang et al. (2024b); Jia et al. (2025), Relevance to the editing intent Paissan et al. (2023); Xu et al. (2024); Zhang et al. (2024b); Manor and Michaeli (2024); Jia et al. (2025); Liu et al. (2024b), and Faithfulness to the original (unedited) content Paissan et al. (2023); Xu et al. (2024); Zhang et al. (2024b); Manor and Michaeli (2024); Jia et al. (2025); Liu et al. (2024b). These criteria provide an intuitive and comprehensive assessment of audio editing performance. Nevertheless, compared to pure generation tasks such as TTA, Text-to-Speech (TTS), or Text-to-Music (TTM), audio editing is inherently more complex, requiring deeper understanding and recognition of nuanced paired contents. This results in increased costs for annotator training, longer evaluation durations, and greater human labor requirements in practice. To these challenges, developing automatic Mean Opinion Score (MOS) prediction methods and frameworks for audio editing is extremely important and meaningful for advancing research in this field.
@@
In this paper, we aim to bridge the data and evaluation gaps in audio editing by introducing AuditEval, a scalable automatic MOS-style assessment framework, and subsequently leveraging it to help mining high-quality pseudo-parallel audio editing data. Our contributions are threefold:
@@
  • Evaluation Dataset Construction. We construct AuditScore, the first comprehensive dataset for subjective audio editing evaluation, consisting of over 6,300 edited pairs from 7+ representative frameworks and 23+ system variants. Each sample is annotated by professional raters from three key aspects: Quality, Relevance, and Faithfulness.
  • Automatic MOS-style Evaluation. Based on AuditScore, we design and systematically study two paradigms for automatic MOS-style assessment in audio editing: a Self-Supervised Learning (SSL)-based evaluator (AuditEval-ssl) and a Large Language Model (LLM)-based evaluator (AuditEval-llm). We aim to provide scalable and cost-effective alternatives to expensive human evaluation.
@@
  • High-Quality Data Mining. We further utilize AuditEval to assess and filter a large amount of synthetically mixed editing pairs, selecting only the most plausible samples to form a high-quality pseudo-parallel subset. Comprehensive experiments validate the effectiveness of this expert-informed filtering strategy in producing superior-quality data, while also highlighting the shortcomings of objective metrics and the advantages of AuditEval in audio editing task.
2 Related Work
2.1 Training-free Audio Editing Method
@@
Training-free audio editing methods exploit pre-trained TTA diffusion models to perform editing without the need for task-specific fine-tuning. These approaches typically begin by inverting the input audio into the noise space of the diffusion model, that is, mapping from $x_{0}$ (the observed audio) to $x_{T}$ (a noisy latent representation). The resulting latent $x_{T}$ preserves the semantic and structural features of the original audio while providing a flexible representation for guided generation. Editing is then achieved by denoising $x_{T}$ into a modified output $\hat{x}_{0}$, conditioned on user instructions or target descriptions. Depending on the specific inversion strategy used, training-free audio editing methods can be broadly categorized into three types: SDEdit-based, DDIM-based, and DDPM-based approaches.
@@
DDIM-based approaches perform deterministic inversion of the diffusion process by exploiting the approximate reversibility of the underlying ODE with sufficiently small step sizes Ho et al. (2020); Song et al. (2020); Mokady et al. (2023). This allows for editing by first reversing the generation trajectory from $\mathbf{z}_{0}$ to $\mathbf{z}T$ (Eq. 2, where $\alpha$ is a hyperparameter and $\varepsilon{\theta}$ denotes the pretrained latent diffusion model), and then regenerating the audio using a modified prompt (Eq. 3), thereby achieving semantically guided edits. Building upon DDIM inversion, several studies  Xu et al. (2024); Zhang et al. (2024b); Jia et al. (2025) incorporate enhancements such as attention manipulation and null-text optimization, achieving effective results in controllable audio editing.
@@
DDPM-based approaches include CycleDiffusion Wu and De la Torre (2023) and Edit-Friendly DDPM Huberman-Spiegelglas et al. (2024). This kind of method leverages the stochastic nature of the DDPM sampling process, specifically, the stochastic noise injected at each diffusion step to preserve the characteristics of the original input. By adding noise to the original input and storing intermediate latent states at each step, the model can compute the corresponding stochastic noise components (Eq.4, $\mu_{t}\left(\mathrm{x}_{t}\right)$ is the calculated expectation of ${x}_{t}$, $\sigma_{t}$ is the hyperparameter), which are then reused during the denoising process under the modified prompt to retain original content. Edit-Friendly DDPM further improves by altering the noise injection mechanism so that the noise added at each timestep is independent, thereby enhancing editing controllability and performance. ZETA Editing Manor and Michaeli (2024) is the first known method to apply DDPM inversion to audio editing, demonstrating superior results compared to SDEdit- and DDIM-inversion-based methods.
@@
2.2 Supervised Audio Editing Models
@@
In contrast to training-free methods, supervised audio editing approaches rely on explicitly constructed editing datasets and can be broadly divided into non-autoregressive and autoregressive paradigms. Models such as Audit and InstructME employ latent diffusion models conditioned on audio-text multimodal inputs. They utilize specialized conditioning mechanisms and position-aware loss functions to enable end-to-end text-controlled audio editing Wang et al. (2023); Han et al. (2023). On the other hand, autoregressive approaches generate edited audio by conditioning on discrete representations of the original audio and textual edit instruction as context. Representative model like MusicGen adopt transformer-based decoders to sequentially synthesize edited audio  Copet et al. (2023); Zhang et al. (2024a). While these models benefit from strong language-audio alignment and high generation quality, they often struggle with temporal inconsistencies between the original and edited audio.
2.3 Auto-Evaluation for Speech, Music and Audio
@@
Automatic subjective evaluation plays a vital role in assessing AI-generated outputs by simulating human perception in a scalable and repeatable manner. Compared to objective metrics, it provides a more intuitive measure of perceptual quality. A well-designed evaluator not only aligns with human preferences but also guides the development of future models. In the speech domain, automatic Mean Opinion Score (MOS) prediction has been widely adopted to evaluate tasks such as Text-to-Speech Cooper et al. (2022); Wang and et al. (2023), Text-to-Voice Tang et al. (2024), and speech enhancement Reddy et al. (2021); Mittag et al. (2021); Chen et al. (2022). These models have shown promising results and effectively complement objective metrics.
@@
In contrast, automatic evaluation for audio and music is more challenging. The content is highly diverse, ranging from natural sound events to complex musical structures. Moreover, the alignment between audio and textual descriptions is inherently looser than the tight correspondence found in speech transcription. To address these challenges, MusicEval Liu et al. (2025) and AudioEval Wang et al. (2025) construct high-quality evaluation datasets and train scoring models to assess both overall quality and text relevance, achieving strong correlations with human judgments. Meanwhile, Meta’s Audiobox Tjandra et al. (2025) proposes a unified framework for speech, music, and audio evaluation, defining four core aesthetic dimensions that collectively provide a comprehensive and structured standard for automatic audio assessment from an aesthetic perspective.
3 Dataset Construction
3.1 Basic Information of AuditScore
Editing Model or Method Generative TTA Backbone Timesteps System IDs Average Editing Scores
Quality Relevance Faithfulness
[…]